OpenStudy (anonymous):

.2 is 5% of what number? solve? Stuck thanks

OpenStudy (anonymous):

\[X \times 5% = 0.2\]\[X \times 5/100 = 0.2\]\[X=100/5 \times 0.5\] X=4

OpenStudy (anonymous):

so we drop the decimal or move it to the left

myininaya (myininaya):

\[.2=5 \% \cdot n\]

OpenStudy (anonymous):

.2=.05*n so divide .2 by .05

OpenStudy (anonymous):

giving me .04

myininaya (myininaya):

\[\%=\frac{1}{100} => .2=5 \cdot \frac{1}{100} \cdot n \text{ solve for n}\]

myininaya (myininaya):

\[.2 \cdot 100 \cdot \frac{1}{5}=n => 20 \cdot \frac{1}{5}=n => n=4\]

OpenStudy (anonymous):

oh ok thank you